Hello,
First off thank you in advance for helping me. My teacher asked me to take this question home and turn it in the day before thanksgiving. This is what we are working on in school. Thanks again.
The slope of a line is m = 2/3. One of the points on the line is (4,-1). Find the equation of the line and present in slope-intercept form.


Why is it useful to write equations in slope-intercept form? What was the form that you used to find the initial equation? What do you need in order to find this equation, could you find an equation with just two points?

Answer by richard1234(7193)

y = mx + b for constants m and b (but we know m = 2/3) (this is the general form for slope-intercept)

y = (2/3)x + b

To find b, plug in x=4, y=-1 to get -1 = (2/3)4 + b, or b = -11/3. Therefore equation: y = (2/3)x - 11/3

Slope-intercept is useful for several purposes, e.g. interpreting linear regressions or how much a dependent variable increases/decreases when the independent variable increases.